OpenStudy (anonymous):

In the decomposition of Th-90, the mass defect = 4.50 x 10^-24 kilograms. If all of this mass is converted to energy, what would be the value in joules? (1 J = 1 kg m2/s2) A) 4.05 x 10^-7 J B) 1.50 x 10^-8 J C) 1.35 x 10^-16 J D) 5.4 x 10^-32 J Can anyone help me figure out how to solve this problem??

OpenStudy (aaronq):

E=mc^2

OpenStudy (anonymous):

so E= 4.50 x 10^-24 / 3.00 x 10^8 ? And then I convert the answer to joules? Or do I have the formula wrong...

OpenStudy (aaronq):

nope, it's multiplied E=(4.50 x 10^-24)(3.00 x 10^8)^2

OpenStudy (anonymous):

So 4.50 x 10^-24 * 3.00 x 10^8 = 1.35 x 10^-15. (1.35 x 10-15)^2 = 1.8225 x 10^-30. Does that look correct? I have to convert that to joules once I get the answer.

OpenStudy (anonymous):

you are only supposed to square the 3.0 x 10^8 NOT the result......that is why you are getting the answer wrong....

OpenStudy (anonymous):

ohhh ok i see

OpenStudy (anonymous):

so do you get the answer??

OpenStudy (anonymous):

I got 0.000000405 :P

OpenStudy (anonymous):

nooo.....did you input the right values on your calculator??

OpenStudy (anonymous):

Wait nevermind I got 1.35 x10^-15

OpenStudy (anonymous):

i did not get the same value...

OpenStudy (anonymous):

I'm confused then :(

OpenStudy (anonymous):

can you type the values you used for the calculation??

OpenStudy (anonymous):

4.50 x 10^-24 * 9.00 x 10^16

OpenStudy (toxicsugar22):

can anyone help me

OpenStudy (anonymous):

if that is exactly what you typed, you should be getting 4.05 x 10^-7

OpenStudy (anonymous):

huh. I don't know what I'm doing wrong then

OpenStudy (anonymous):

i've tried the same calculation a dozen times and still get the same answer :/

OpenStudy (anonymous):

oh alright I guess I was just going wrong somewhere. So now I convert that joules.

OpenStudy (anonymous):

no..that is it...its already on joules....

OpenStudy (anonymous):

ohhh okay. Well thank you for helping :)

OpenStudy (anonymous):

you are welcome... if you are given a question next time with the mass in grams, you would have to convert it into kilograms first....then after that your energy calculated would be in joules..

OpenStudy (anonymous):

Alright that makes sense. Thanks